OpenStudy (anonymous):

how do i solve 3x+2y=12 x+4y=14 Using the elimination method

OpenStudy (anonymous):

multiply the 2nd equation with 3 and subtract from the first one

OpenStudy (anonymous):

You can multiply the first equation with 2. You'd then get 6x+4y=24. Subtract the second equation from it and you have "eliminated y". Solve for x, and then for y using the value of x.

OpenStudy (anonymous):

To solve for these equations pick any equation solve for y or x lets solve for x in the second equation so we got x = 14 - 4y now let us sub this into the first equation and we got 3(14-4y) +2y = 12 now we just solve for y which is 42 - 12y +2y =12 and by collecting like terms 30=10y so y=3 and now let us sub this into our first equation of x so x=14 - 4(3) = 2 so x=2 and y=3 :)